OpenStudy (anonymous):

Solve 2 cos^2 y - sin y - 1 = 0, for 0 ≤ y ≤360 degrees.

OpenStudy (anonymous):

The values don't work.

OpenStudy (anonymous):

2(1-sin^2y)-siny-1=0 1-siny-2sin^2y=0 (2siny-1)(siny+1)=0 siny=-1 or 1/2 y=270,30,150

OpenStudy (anonymous):

why??

OpenStudy (anonymous):

when y=270 then cosy=0, siny=-1 s0 0-(-1)-1=0

OpenStudy (anonymous):

change ur calculator mode from radian to degree...or put the eqivalent radians..of my ans.

OpenStudy (anonymous):

see i have shown that y=270 is a root...similarly u can do it by urself for 30 and 150

OpenStudy (anonymous):

Sorry, I just realized you're right. Thanks a lot! :)

OpenStudy (anonymous):

good to hear that u have got it..